You are starting a T-shirt business. your fixed costs are $5,000. It costs $11 to buy each shirt. You plan to sell the shirts for $16 each. Write the total cost function, revenue function, profit function. Find the break even point.
Answer by checkley77(12844)

16-11=$5 is the per shirt profit.
5,000+11x=16x
11x-16x=-5,000
-5x=-5,000
x=-5,000/-5
x=1,000 shirts need to be sold to break even.
revenue=16*1,000=$16,000 is the revenue figure.
cost=5,000+1,000+11,000=$16,000 is the cost factor.